# Moving through Cartesian products, coronas and joins in general position

### v1

@@ Abstract
e Department of Informatics and Statistics, Klaipėda University, Lithuania
f Departamento de Estadística e IO, Universidad de Cádiz, Algeciras Campus, Spain
@@ 1 Introduction
However, this picture lacks an important feature of real world robotic navigation: the general position problem is ‘static’, whereas in applications the robots will typically need to move around the network. Watching the mobile delivery robots created by Starship Technologies® [35] inspired the authors of [16] to consider a dynamic version of the general position problem, in which robots move through the vertices of a graph whilst remaining in general position, and such that every vertex is visited by a robot at least once (it is assumed that one robot moves to an adjacent vertex at each step and no vertex can contain more than one robot). The paper [16] considered this problem for block graphs, rooted products, Kneser graphs, unicyclic graphs, complete multipartite graphs and line graphs of complete graphs. A mobile version of the closely related mutual-visibility problem was recently treated in [10]. In addition, the paper [4] considers the version of the mobile general position problem with the stricter condition that every vertex must be visited by every robot, as well as scenarios in which more than one robot is allowed to move at once.
The general position problem has been investigated for a wide variety of graphs, but there is a particularly extensive literature on general position sets in Cartesian products, see [17, 18, 20, 32, 33]. General position sets in other graph products were discussed in [14]. Cartesian products have also been considered in the setting of variants of the general position number, such as the lower general position number [21], the mutual-visibility number [9, 12], the monophonic position number [7], the lower mutual-visibility number [5], general position polynomials [15], edge general position numbers [27], total mutual-visibility [23] and the variety of general position problems [31]. In this paper, we examine the mobile general position problem in Cartesian products, together with the coronas and joins of graphs.
@@ 1.1 Formal definitions and preliminaries
A path $P_{r}$ in $G$ is a sequence $u_{1},u_{2},\dots,u_{r}$ such that $u_{i}\sim u_{i+1}$ for $i\in[r-1]$ and the length of this path is $r-1$. A cycle $C_{r}$ is a sequence $u_{1},u_{2},\dots,u_{r}$ such that $u_{i}\sim u_{i+1}$ for $i\in[r-1]$ and also $u_{1}\sim u_{r}$. The distance between vertices $u,v\in V(G)$ is the length of a shortest $u,v$-path in $G$. We will typically identify the vertices of a cycle $C_{n}$ with $\mathbb{Z}_{n}$ and the vertices of a path $P_{n}$ with $[n]$ in the natural manner. A subgraph $X$ of $G$ is convex if for any $u,v\in V(X)$ we have $V(P)\subseteq V(X)$ for any shortest $u,v$-path $P$ in $G$. The subgraph $G[X]$ induced by a subset $X\subseteq V(G)$ is the subgraph with vertex set $X$ such that $u,v\in X$ are adjacent in $G[X]$ if and only if they are adjacent in $G$.
If $S\subseteq V(G)$ has the property that no shortest path of $G$ passes through more than two vertices of $S$, then we say that $S$ is a general position set of $G$. The largest possible number of vertices in a general position set of $G$ is the general position number of $G$, denoted by $\operatorname{gp}(G)$. Any largest general position set is referred to as a gp-set.
@@
If a robot is located at a vertex $u$ and $u\sim v$, then we indicate the movement of the robot from $u$ to $v$ along the edge $uv$ by $u\rightsquigarrow v$ and refer to this as a move. Suppose that we assign exactly one robot to each vertex of a general position set $S$. If a robot is stationed at a vertex $u$ of $S$, then the move $u\rightsquigarrow v$ is called a legal move if i) $v\notin S$ (thereby avoiding having more than one robot per vertex at any stage) and ii) the new set $(S\setminus\{u\})\cup\{v\}$ is also a general position set. A configuration of robots on a general position set of $G$ is called a mobile general position set if there is a sequence of legal moves starting from $S$ such that every vertex of $G$ is visited at least once by some robot. The mobile general position number, written $\operatorname{Mob_{gp}}(G)$, is the largest number of robots in a mobile general position set. We will refer to a largest possible configuration of robots in mobile general position as a $\operatorname{Mob_{gp}}$-set.
@@
In [31] the variety of general position problems in graphs was introduced, including general position sets, outer general position sets, dual general position sets, and total general position sets. For our purposes we recall the following definitions. If $X\subseteq V(G)$, then $u,v\in V(G)$ are $X$-positionable if for any shortest $u,v$-path $P$ we have $V(P)\cap X\subseteq\{u,v\}$. Note that then $X$ is a general position set if all pairs $u,v\in X$ are $X$-positionable. If it also holds that every pair $u,v$ with $u\in X$ and $v\in V(G)\setminus X$ is $X$-positionable, then $X$ is an outer general position set. The cardinality of a largest outer general position set of $G$ is denoted by $\operatorname{gp}_{\rm o}(G)$ and is called the outer general position number. It is shown in [31] that outer general position sets coincide with sets of mutually maximally distant vertices. (The latter concept was introduced in [28], see also the related survey [22].) In particular, in a block graph the outer general position sets are the sets of simplicial vertices; in the case of a tree this yields $\operatorname{gp}(T)=\operatorname{gp}_{\rm o}(T)=\ell(T)$.
@@ 2 Bounds for Cartesian products
$(i)$ Let $S$ be a $\operatorname{Mob_{gp}}$-set of $G$ and let $h\in V(H)$. We first position the robots at the vertices of the set $S\times\{h\}$. As $G^{h}$ is an isometric subgraph of $G\operatorname{\,\square\,}H$, robots initially stationed at the vertices of $S\times\{h\}$ can visit every vertex of $G^{h}$ by a sequence of legal moves, all the time remaining inside $G^{h}$ and in in general position in $G\operatorname{\,\square\,}H$. Now, whenever a robot visits a vertex $(g,h)$ in this layer, this robot can visit all the vertices in the $H$-layer ${}^{g}H$ corresponding to $g$ by a sequence of legal moves and then return to $(g,h)$. Hence $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}H)\geq\operatorname{Mob_{gp}}(G)$. By symmetry, $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}H)\geq\operatorname{Mob_{gp}}(H)$.
@@
$(ii)$ Let $S=\{u_{1},\dots,u_{r}\}$ be an outer general position set of $G$ of cardinality $\operatorname{gp}_{\rm o}(G)$ and start with that many robots positioned on $S\times\{h\}$ for some $h\in V(H)$. Let $R_{i}$ be the robot at $(u_{i},h)$ for $i\in[r]$. Also, for each $i\in[r]$ let $G_{i}$ be the connected component containing $u_{i}$ in $G-(S\setminus\{u_{i}\})$.
@@
Let $h^{\prime}\in V(H)\setminus\{h\}$ and let $Q$ be a shortest $h,h^{\prime}$-path in $H$. The robot $R_{i}$ can follow the path ${}^{u_{i}}Q$ from $(u_{i},h)$ to reach the vertex $(u_{i},h^{\prime})$ by legal moves. At this point, $R_{i}$ can visit all the vertices of $V(G_{i})\times\{h^{\prime}\}$ by a sequence of legal moves. To see this, notice that the shortest paths between the remaining robots lie within the layer $G^{h}$, whilst the shortest paths from $R_{i}$ to any $R_{j}$, $i\neq j$, do not pass through a third robot by the outer general position property. Afterwards $R_{i}$ can return to $(u_{i},h)$ by performing these legal moves in the reverse order. As this holds for any of the robots and any $h^{\prime}\in V(H)-\{h\}$, this allows us to perform a sequence of legal move so that any vertex of $G\operatorname{\,\square\,}H$ outside $G^{h}$ is visited, since $\bigcup_{i=1}^{r}V(G_{i})=V(G)$. Having done this, we return all the robots to their original positions in $S\times\{h\}$ using legal moves.
@@
Finally, let $h^{\prime}\in N_{H}(h)$. For $i\in[r]$ we move the robot $R_{i}$ from $(u_{i},h)$ to $(u_{i},h^{\prime})$ in sequence; as $S$ is in general position each of these moves is legal. Afterwards the robots will occupy the set $S\times\{h^{\prime}\}$. The previous reasoning applied to $G^{h^{\prime}}$ now shows that the robots visit each vertex of $V(G)\times\{h\}$ by legal moves. Thus $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}H)\geq\operatorname{gp}_{\rm o}(G)$ and, by a symmetric argument, $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}H)\geq\operatorname{gp}_{\rm o}(H)$. ∎
@@
We now show that both lower bounds in Proposition 2.1 are sharp by considering the Cartesian products $K_{r}\operatorname{\,\square\,}P_{s}$. We recall that [32, Theorem 3.2] implies that $\operatorname{gp}(K_{r}\operatorname{\,\square\,}P_{s})=r+1$ for $s\geq 3$. Observe that for $r,s\geq 2$ it holds that $\max\{\operatorname{Mob_{gp}}{(K_{r})},\operatorname{Mob_{gp}}{(P_{s})}\}=\max\{\operatorname{gp}_{\rm o}{(K_{r})},\operatorname{gp}_{\rm o}{(P_{s})}\}=r$.
@@
Let $r,s\geq 2$. By Proposition 2.1 we have $\operatorname{Mob_{gp}}(K_{r}\operatorname{\,\square\,}P_{s})\geq r$. We now show that $\operatorname{Mob_{gp}}(K_{r}\operatorname{\,\square\,}P_{s})\leq r$. Let $V(P_{s})=[s]$ and suppose for a contradiction that there exists a mobile general position set $S$ of $K_{r}\operatorname{\,\square\,}P_{s}$ with $|S|>r$. Then choose $x\in V(K_{r})$, such that $(x,i),(x,j)\in V(K_{r}\operatorname{\,\square\,}P_{s})$ are occupied by robots $R_{1},R_{2}$, where $i<j$ and the difference $|j-i|$ is the smallest possible.
First, notice that any other vertex $(y,k)$ with $k\geq j$ cannot be occupied by a robot of $S$, for otherwise $(x,j)$ would lie on a shortest path between $(x,i)$ and $(y,k)$. Similarly, every other vertex $(y,k)$ with $k\leq i$ is unoccupied. If $s=2$, this leads us to conclude that $|S|\leq 2$, which is impossible, since we assumed that $|S|>r\geq 2$. So, we may now suppose that $s>2$. By the above reasoning, and the fact that the difference $|j-i|$ was chosen to be as small as possible, we deduce that no layer ${}^{y}P_{s}$ can contain more than one robot for $y\in V(K_{r})\setminus\{x\}$. Thus, we must have $|S|=r+1$ and each layer ${}^{y}P_{s}$ contains exactly one robot for $y\neq x$. Moreover, each of the remaining $r-1$ robots different from $R_{1}$ and $R_{2}$ are located at vertices $(y,\ell)$ such that $i<\ell<j$.
It now follows that neither $R_{1}$ nor $R_{2}$ can cross to another $P_{s}$-layer by a legal move. Suppose that robots $R_{1}$ and $R_{2}$ are stationed at $(x,i)$ and $(x,j)$, respectively, when $R_{1}$ makes the move $(x,i)\rightsquigarrow(y,i)$. Since there is a robot at some vertex $(y,k)$ with $i<k<j$, after this move the robot at $(y,k)$ would lie on a shortest path from $(y,i)$ to $(x,j)$. As a result, no robot can visit any vertex in $(V(K_{n})\setminus\{x\})\times\{1,s\}$, a contradiction. Thus at most $r$ robots can traverse $K_{r}\operatorname{\,\square\,}P_{s}$ in general position. ∎
To see that the lower bounds of Proposition 2.1 are independent in general, consider the following examples. If $n\geq 2$, then $\operatorname{Mob_{gp}}(K_{n,n})=2$ and $\operatorname{gp}_{\rm o}(K_{n,n})=n$. Hence the bound $(i)$ yields $\operatorname{Mob_{gp}}(K_{n,n}\operatorname{\,\square\,}K_{n,n})\geq 2$, while $(ii)$ yields $\operatorname{Mob_{gp}}(K_{n,n}\operatorname{\,\square\,}K_{n,n})\geq n$. On the other hand, if $n\geq 7$, then $\operatorname{Mob_{gp}}(C_{n})=3$ and $\operatorname{gp}_{\rm o}(C_{n})=2$, hence the bound $(i)$ is better for $\operatorname{Mob_{gp}}(C_{n}\operatorname{\,\square\,}C_{n})$ if $n\geq 7$. Moreover, just after Theorem 2.4 we will demonstrate that the mobile general position number of a graph can be arbitrarily larger than its outer general position number, so that in turn, bound $(i)$ can be arbitrarily larger than bound $(ii)$.
@@
Let $X$ be a gp-set of $K_{n}\operatorname{\,\square\,}K_{m}$. Then, as stated above, $|X|=n+m-2$. Note that the set $X$ contains at most two vertices from every induced copy of $C_{4}$ of $K_{n}\operatorname{\,\square\,}K_{m}$. This fact implies the following:
@@
If $X$ contains all the vertices of some $K_{n}$-layer, then by Claim A, $X$ contains no other vertices, implying that $|X|=n<n+m-2$, which is not possible. Let $j\in[m]$ be such that $t=|X\cap V((K_{n})^{j})|$ is as large as possible. Note that $t\geq 2$, for otherwise we would have $|X|\leq m$. Moreover, by the above, $t\leq n-1$. If $t=n-1$ and $i\in[n]$ is the index for which $(i,j)\notin X$, then using Claim A again, we have that $X\subseteq V((K_{n})^{j})\cup V(^{i}(K_{m}))$. Since $|X|=n+m-2$, we conclude that $X$ has the required structure in this case. Suppose finally that $t=n-k\geq 2$, where $k\geq 2$. We may assume without loss of generality that $X\cap V((K_{n})^{j})=\{(1,j),\ldots,(n-k,j)\}$. Then, using Claim A once more,
@@
We have thus proved that $|X|=n+m-2$ holds only in the case when $t=n-1$ and $X$ has the structure as claimed. ∎
Lemma 2.3 is illustrated in Fig. 1, where the gp-set of $K_{7}\operatorname{\,\square\,}K_{5}$ corresponding to the vertex $(i,j)$ is shown.
$i$$j$
@@
Let $m\geq 3$. By Lemma 2.3, every gp-set of $K_{n}\operatorname{\,\square\,}K_{m}$ has the canonical form as illustrated in Fig. 1. It is straightforward to check that no robot placed in such a set can make a legal move. This implies that $\operatorname{Mob_{gp}}(K_{n}\operatorname{\,\square\,}K_{m})<\operatorname{gp}(K_{n}\operatorname{\,\square\,}K_{m})=n+m-2$. To complete the proof, we are going to show that there exists a mobile general position set of cardinality $n+m-3$.
Let $X=\{(2,1),(3,1),\ldots,(n,1)\}\cup\{(1,3),(1,4),\ldots,(1,m)\}$ and let us first position a set of $n+m-3$ robots at the vertices of $S$. First notice that such a set is a general position set, since it is a subset of the canonical gp-set of $K_{n}\operatorname{\,\square\,}K_{m}$, as illustrated in Fig. 1. Then we perform the following sequence of moves:
@@
Observe that each of these moves is legal. Thus, the new set occupied by the robots is a general position set of $K_{n}\operatorname{\,\square\,}K_{m}$. Next, this process of legal moves can be repeated $m-2$ times. In this way, the robots initially positioned at $(2,1),(3,1),\ldots,(n,1)$ will visit the vertices of the set $\{2,3,\ldots,n\}\times[m]$ by legal moves, while the remaining vertices can be visited by legal moves by the robots initially positioned at $(1,3),(1,4),\ldots,(1,m)$. This concludes our argument for the existence of a mobile general position set of cardinality $n+m-3$. ∎
By Theorem 2.4 we have $\operatorname{Mob_{gp}}(K_{n}\operatorname{\,\square\,}K_{n})=2n-3$, whilst it follows from [31, Theorem 2.3] that $\operatorname{gp}_{\rm o}(K_{n}\operatorname{\,\square\,}K_{n})=n$, so that the mobile general position number of a graph can be arbitrarily larger than the outer general position number. Theorem 2.4 also demonstrates that the mobile general position number of a Cartesian product can be arbitrarily larger than both bounds in Proposition 2.1.

### v2

@@ Abstract
e Departamento de Estadística e IO, Universidad de Cádiz, Algeciras Campus, Spain
f Department of Informatics and Statistics, Klaipėda University, Lithuania
@@ 1 Introduction
However, this picture lacks an important feature of real world robotic navigation: the general position problem is ‘static’, whereas in applications the robots will typically need to move around the network. Watching the mobile delivery robots created by Starship Technologies® [32] inspired the authors of [17] to consider a dynamic version of the general position problem, in which robots move through the vertices of a graph whilst remaining in general position, and such that every vertex is visited by a robot at least once (it is assumed that one robot moves to an adjacent vertex at each step and no vertex can contain more than one robot). The paper [17] considered this problem for block graphs, rooted products, Kneser graphs, unicyclic graphs, complete multipartite graphs and line graphs of complete graphs. A mobile version of the closely related mutual-visibility problem was recently treated in [10]. In addition, the paper [3] considers the version of the mobile general position problem with the stricter condition that every vertex must be visited by every robot.
The general position problem has been investigated for a wide variety of graphs, but there is a particularly extensive literature on general position sets in Cartesian products, see [18, 19, 20, 30, 31]. General position sets in other graph products were discussed in [14]. Cartesian products have also been considered in the setting of variants of the general position number, such as the lower general position number [21], the mutual-visibility number [9, 12], the monophonic position number [7], the lower mutual-visibility number [4], general position polynomials [15], edge general position numbers [27], total mutual-visibility [23], the variety of general position problems [29] and general position and mutual-visibility colourings [5, 16]. In this paper, we examine the mobile general position problem in Cartesian products, together with the coronas and joins of graphs.
@@ 1.1 Formal definitions and preliminaries
A path $P_{r}$ in $G$ is a sequence $u_{1},u_{2},\dots,u_{r}$ of distinct vertices such that $u_{i}\sim u_{i+1}$ for $i\in[r-1]$ and the length of this path is r1r-1. A cycle $C_{r}$ is a sequence $u_{1},u_{2},\dots,u_{r}$ such that $u_{i}\sim u_{i+1}$ for $i\in[r-1]$ and also $u_{1}\sim u_{r}$. The distance between vertices $u,v\in V(G)$ is the length of a shortest $u,v$-path in $G$. We will typically identify the vertices of a cycle $C_{n}$ with $\mathbb{Z}_{n}$ and the vertices of a path $P_{n}$ with $[n]$ in the natural manner. A subgraph $X$ of $G$ is convex if for any $u,v\in V(X)$ we have $V(P)\subseteq V(X)$ for any shortest $u,v$-path $P$ in $G$. The subgraph $G[X]$ induced by a subset $X\subseteq V(G)$ is the subgraph with vertex set $X$ such that $u,v\in X$ are adjacent in $G[X]$ if and only if they are adjacent in $G$.
If $S\subseteq V(G)$ has the property that no three vertices of $S$ lie on a common shortest path of $G$, then we say that $S$ is a general position set of $G$. The largest possible number of vertices in a general position set of $G$ is the general position number of $G$, denoted by $\operatorname{gp}(G)$. Any largest general position set is referred to as a gp-set. Fig. 1 shows two general position sets in the well-known Petersen graph. The red vertices on the left form a maximal but non-optimal general position set of order four, whilst the six red vertices on the right constitute a $\operatorname{gp}$-set of the Petersen graph, and so the Petersen graph has general position number six. Note that in both cases no shortest path between any pair of red vertices passes through a third red vertex.
Figure 1: A (non-optimal) general position set (left) and a gp-set (right) of the Petersen graph are shown in red.
@@
If a robot is located at a vertex $u$ and $u\sim v$, then we indicate the movement of the robot from $u$ to $v$ along the edge $uv$ by $u\rightsquigarrow v$ and refer to this as a move. Suppose that we assign exactly one robot to each vertex of a general position set $S$. If a robot is stationed at a vertex $u$ of $S$, then the move $u\rightsquigarrow v$ is called a legal move if i) $v\notin S$ (thereby avoiding having more than one robot per vertex at any stage) and ii) the new set $(S\setminus\{u\})\cup\{v\}$ is also a general position set. A configuration of robots on a general position set of $G$ is called a mobile general position set if there is a sequence of legal moves starting from $S$ such that every vertex of $G$ is visited at least once by some robot. The mobile general position number, written $\operatorname{Mob_{gp}}(G)$, is the largest number of robots in a mobile general position set of $G$. We will refer to a largest possible configuration of robots in mobile general position as a $\operatorname{Mob_{gp}}$-set.
Fig. 2 shows a few configurations of a $\operatorname{Mob_{gp}}$-set for the Petersen graph. The red arrows indicate the legal moves that lead from one configuration to the next. Notice that the last configuration is equivalent to the first up to symmetry. By repeating (symmetrical equivalents of) this sequence of legal moves every vertex of the Petersen graph can be visited. It is shown in [17] that this mobile general position has the largest possible cardinality amongst all such sets, i.e. it is a $\operatorname{Mob_{gp}}$-set.
Figure 2: A sequence of legal moves (shown by red arrows) for a $\operatorname{Mob_{gp}}$-set of the Petersen graph.
@@
In [29] the variety of general position problems in graphs was introduced, including general position sets, outer general position sets, dual general position sets, and total general position sets. For our purposes we recall the following definitions. If $X\subseteq V(G)$, then $u,v\in V(G)$ are $X$-positionable if for any shortest $u,v$-path $P$ we have $V(P)\cap X\subseteq\{u,v\}$. Hence $X$ is a general position set if all pairs $u,v\in X$ are $X$-positionable. If it also holds that every pair $u,v$ with $u\in X$ and $v\in V(G)\setminus X$ is $X$-positionable, then $X$ is an outer general position set. The cardinality of a largest outer general position set of $G$ is denoted by $\operatorname{gp}_{\rm o}(G)$ and is called the outer general position number. It is shown in [29] that outer general position sets coincide with sets of mutually maximally distant vertices. (The latter concept was introduced in [28], see also the related survey [22].) In particular, in a block graph the outer general position sets are the sets of simplicial vertices; in the case of a tree this yields $\operatorname{gp}(T)=\operatorname{gp}_{\rm o}(T)=\ell(T)$.
@@ 2 Bounds for Cartesian products
$(i)$ Let $S$ be a $\operatorname{Mob_{gp}}$-set of $G$ and let $h\in V(H)$. We first position the robots at the vertices of the set $S\times\{h\}$. As $G^{h}$ is a convex subgraph of $G\operatorname{\,\square\,}H$, robots initially stationed at the vertices of $S\times\{h\}$ can visit every vertex of $G^{h}$ by a sequence of legal moves, all the time remaining inside $G^{h}$ and in general position in $G\operatorname{\,\square\,}H$. Now, whenever a robot visits a vertex $(g,h)$ in this layer, this robot can visit all the vertices in the $H$-layer $^{g}H$ corresponding to $g$ by a sequence of legal moves and then return to $(g,h)$. Hence $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}H)\geq\operatorname{Mob_{gp}}(G)$. By symmetry, $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}H)\geq\operatorname{Mob_{gp}}(H)$.
@@
$(ii)$ Let $S=\{u_{1},\dots,u_{r}\}$ be an outer general position set of $G$ of cardinality $\operatorname{gp}_{\rm o}(G)$ and start with robots positioned at each vertex of $S\times\{h\}$ for some $h\in V(H)$. Let $R_{i}$ be the robot at $(u_{i},h)$ for $i\in[r]$. Also, for each $i\in[r]$, let $G_{i}$ be the connected component containing $u_{i}$ in $G\setminus(S\setminus\{u_{i}\})$.
@@
Let $h^{\prime}\in V(H)\setminus\{h\}$ and let $Q$ be a shortest $h,h^{\prime}$-path in $H$. The robot $R_{i}$ can follow the path $^{u_{i}}Q$ from $(u_{i},h)$ to reach the vertex $(u_{i},h^{\prime})$ by legal moves. At this point, $R_{i}$ can visit all the vertices of $V(G_{i})\times\{h^{\prime}\}$ by a sequence of legal moves. To see this, notice that the shortest paths between the robots remaining in $G^{h}$ lie within the layer $G^{h}$, whilst the shortest paths from $R_{i}$ to any $R_{j}$, $i\neq j$, do not pass through a third robot by the outer general position property. Afterwards $R_{i}$ can return to $(u_{i},h)$ by performing these legal moves in the reverse order. As this holds for any of the robots and any $h^{\prime}\in V(H)\setminus\{h\}$, this allows us to perform a sequence of legal moves so that any vertex of $G\operatorname{\,\square\,}H$ outside $G^{h}$ is visited, since $\bigcup_{i=1}^{r}V(G_{i})=V(G)$. Having done this, we return all the robots to their original positions in $S\times\{h\}$ using legal moves.
@@
Finally, let $h^{\prime}\in N_{H}(h)$. For $i\in[r]$ we move the robot $R_{i}$ from $(u_{i},h)$ to $(u_{i},h^{\prime})$ in sequence; as $S$ is in general position each of these moves is legal. Afterwards the robots will occupy the set $S\times\{h^{\prime}\}$. The previous reasoning applied to $G^{h^{\prime}}$ now shows that the robots can visit each vertex of $V(G)\times\{h\}$ by legal moves. Thus $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}H)\geq\operatorname{gp}_{\rm o}(G)$ and, by a symmetric argument, $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}H)\geq\operatorname{gp}_{\rm o}(H)$. ∎
@@
We now show that both lower bounds in Proposition 2.1 are sharp by considering the Cartesian products $K_{r}\operatorname{\,\square\,}P_{s}$. We recall that [30, Theorem 3.2] implies that $\operatorname{gp}(K_{r}\operatorname{\,\square\,}P_{s})=r+1$ for $s\geq 3$. Observe that for $r,s\geq 2$ it holds that $\max\{\operatorname{Mob_{gp}}{(K_{r})},\operatorname{Mob_{gp}}{(P_{s})}\}=\max\{\operatorname{gp}_{\rm o}{(K_{r})},\operatorname{gp}_{\rm o}{(P_{s})}\}=r$.
@@
Let $r,s\geq 2$. By Proposition 2.1 we have $\operatorname{Mob_{gp}}(K_{r}\operatorname{\,\square\,}P_{s})\geq r$. We now show that $\operatorname{Mob_{gp}}(K_{r}\operatorname{\,\square\,}P_{s})\leq r$. Let $V(P_{s})=[s]$ and suppose for a contradiction that there exists a mobile general position set $S$ of $K_{r}\operatorname{\,\square\,}P_{s}$ with $|S|>r$. Then choose $x\in V(K_{r})$, such that $(x,i),(x,j)\in V(K_{r}\operatorname{\,\square\,}P_{s})$ are occupied by robots $R_{1},R_{2}$, where $i<j$.
First, notice that for any $y\in V(K_{r})$ and $j\leq k\leq s$ the vertex $(y,k)$ cannot be occupied by a robot in the initial configuration $S$, for otherwise $(x,j)$ would lie on a shortest path between $(x,i)$ and $(y,k)$. Similarly, every other vertex $(y,k)$ with $k\leq i$ is unoccupied. The same reasoning shows that any other $P_{s}$-layer can contain at most one robot, since if there are robots at $(y,k)$ and $(y,k^{\prime})$, where $y\in V(K_{r})\setminus\{x\}$ and $i<k<k^{\prime}<j$, then $(y,k^{\prime})$ would lie on a shortest $(y,k),(x,j)$-path. Thus, we must have $|S|=r+1$ and each layer $^{y}P_{s}$ contains exactly one robot for $y\neq x$. Moreover, each of the remaining r1r-1 robots different from $R_{1}$ and $R_{2}$ are located at vertices $(y,k)$ such that $i<k<j$.
It now follows that neither $R_{1}$ nor $R_{2}$ can cross to another $P_{s}$-layer by a sequence of legal moves. Otherwise, suppose that robots $R_{1}$ and $R_{2}$ are stationed at $(x,i^{\prime})$ and $(x,j^{\prime})$ respectively just before $R_{1}$ moves to a different $P_{s}$-layer by the legal move $(x,i^{\prime})\rightsquigarrow(y,i^{\prime})$, $y\in V(K_{r})\setminus\{x\}$. Since there is a robot at some vertex $(y,k^{\prime})$ with $i^{\prime}<k^{\prime}<j^{\prime}$, after this move the robot at $(y,k^{\prime})$ would lie on a shortest path from $(y,i^{\prime})$ to $(x,j^{\prime})$. As a result, no robot can visit any vertex in $(V(K_{n})\setminus\{x\})\times\{1,s\}$, a contradiction. Thus at most $r$ robots can traverse $K_{r}\operatorname{\,\square\,}P_{s}$ in general position. ∎
To see that the lower bounds of Proposition 2.1 are independent in general, consider the following examples. If $n\geq 2$, then $\operatorname{Mob_{gp}}(K_{n,n})=2$ and $\operatorname{gp}_{\rm o}(K_{n,n})=n$. Hence the bound $(i)$ yields $\operatorname{Mob_{gp}}(K_{n,n}\operatorname{\,\square\,}K_{n,n})\geq 2$, whilst $(ii)$ yields $\operatorname{Mob_{gp}}(K_{n,n}\operatorname{\,\square\,}K_{n,n})\geq n$. On the other hand, if $n\geq 7$, then $\operatorname{Mob_{gp}}(C_{n})=3$ and $\operatorname{gp}_{\rm o}(C_{n})=2$, hence the bound $(i)$ is better for $\operatorname{Mob_{gp}}(C_{n}\operatorname{\,\square\,}C_{n})$ if $n\geq 7$. Moreover, just after Theorem 2.4 we will demonstrate that the mobile general position number of a graph can be arbitrarily larger than its outer general position number, so that bound $(i)$ can also be arbitrarily larger than bound $(ii)$.
@@
Let $X$ be a gp-set of $K_{n}\operatorname{\,\square\,}K_{m}$. Then, as stated above, |X|=n+m2|X|=n+m-2. Note that the set $X$ contains at most two vertices from every induced copy of $C_{4}$ of $K_{n}\operatorname{\,\square\,}K_{m}$. This fact implies the following:
@@
If $X$ contains all the vertices of some $K_{n}$-layer, then by Claim A, $X$ contains no other vertices, implying that |X|=n<n+m2|X|=n<n+m-2, which is not possible. Let $j\in[m]$ be such that $t=|X\cap V((K_{n})^{j})|$ is as large as possible. Note that $t\geq 2$, for otherwise we would have $|X|\leq m$. Moreover, by the above, $t\leq n-1$. If t=n1t=n-1 and $i\in[n]$ is the index for which $(i,j)\notin X$, then using Claim A again, we have that $X\subseteq V((K_{n})^{j})\cup V(^{i}(K_{m}))$. Since |X|=n+m2|X|=n+m-2, we conclude that $X$ has the required structure in this case. Suppose finally that $t=n-k\geq 2$, where $k\geq 2$. We may assume without loss of generality that $X\cap V((K_{n})^{j})=\{(1,j),\ldots,(n-k,j)\}$. Then, using Claim A once more,
@@
We have thus proved that |X|=n+m2|X|=n+m-2 holds only in the case when t=n1t=n-1 and $X$ has the structure as claimed. ∎
Lemma 2.3 is illustrated in Fig. 3, where the gp-set of $K_{7}\operatorname{\,\square\,}K_{5}$ corresponding to the vertex $(i,j)$ is shown. We now use this result to find the mobile general position number of $K_{n}\operatorname{\,\square\,}K_{m}$.
iijj
@@
Let $m\geq 3$. By Lemma 2.3, every gp-set of $K_{n}\operatorname{\,\square\,}K_{m}$ has the canonical form as illustrated in Fig. 3. It is straightforward to check that no robot placed in such a set can make a legal move. This implies that $\operatorname{Mob_{gp}}(K_{n}\operatorname{\,\square\,}K_{m})<\operatorname{gp}(K_{n}\operatorname{\,\square\,}K_{m})=n+m-2$.
To complete the proof, we are going to show that there exists a mobile general position set of cardinality n+m3n+m-3. Station n+m3n+m-3 robots on the vertices of $S=\{(2,1),(3,1),\ldots,(n,1)\}\cup\{(1,3),(1,4),\ldots,(1,m)\}$. This set is a general position set, since it is a subset of the canonical gp-set of $K_{n}\operatorname{\,\square\,}K_{m}$, as illustrated in Fig. 3. Then we perform the following sequence of moves:
@@
Observe that each of these moves is legal. Thus, the new set occupied by the robots is a general position set of $K_{n}\operatorname{\,\square\,}K_{m}$. Next, this process of legal moves can be repeated m2m-2 times, i.e. for $2\leq j\leq m-1$ perform the sequence
$(2,j)\rightsquigarrow(2,j+1),(3,j)\rightsquigarrow(3,j+1),\ldots,(n,j)\rightsquigarrow(n,j+1),(1,j+2)\rightsquigarrow(1,j-1),$
(skipping the final undefined move). In this way, the robots initially positioned at $(2,1),(3,1),\ldots,(n,1)$ will visit the vertices of the set $\{2,3,\ldots,n\}\times[m]$ by legal moves, while the remaining vertices can be visited by legal moves by the robots initially positioned at $(1,3),(1,4),\ldots,(1,m)$. This concludes our argument for the existence of a mobile general position set of cardinality n+m3n+m-3. ∎
By Theorem 2.4 we have $\operatorname{Mob_{gp}}(K_{n}\operatorname{\,\square\,}K_{n})=2n-3$, whilst it follows from [29, Corollary 4.4] that $\operatorname{gp}_{\rm o}(K_{n}\operatorname{\,\square\,}K_{n})=n$, so that the mobile general position number of a graph can be arbitrarily larger than the outer general position number. Theorem 2.4 also demonstrates that the mobile general position number of a Cartesian product can be arbitrarily larger than both bounds in Proposition 2.1.